OpenStudy (anonymous):

sec x + 1 = tan x? How do I solve?

OpenStudy (raden):

first, squares both sides. what do you get ...

OpenStudy (anonymous):

sec^2 x+1=tan^2 x

OpenStudy (raden):

sec^2 x + 2 secx + 1 = tan^2 x remember that ; (a+b)^2 = a^2 + 2ab + b^2